OpenStudy (anonymous):

A 12 foot ladder is leaning against a wall. If the top slips down the wall at a rate of 2 ft/s, how fast will the foot be moving away from the wall when the top is 11 feet above the ground? The foot will be moving at__________ft/s.

OpenStudy (anonymous):

3.33 ft/s

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

any time

OpenStudy (anonymous):

wait i got it wrong,

OpenStudy (anonymous):

what ?

OpenStudy (anonymous):

can you give me some more decimals?

OpenStudy (anonymous):

look the rational expression is \[\frac{ 20 }{ \sqrt{44} }\]

OpenStudy (anonymous):

im still getting it wrong :(

OpenStudy (anonymous):



OpenStudy (anonymous):

but excuse me i plugged h for 10 instead of 11 sorry !!

OpenStudy (anonymous):

just a sec

OpenStudy (anonymous):

4.5929

OpenStudy (anonymous):

im still getting it wrong for some reason, i just cant figure it out

OpenStudy (anonymous):

whaaaaa ?? do you know how to solve it ??

OpenStudy (anonymous):

like I got the same thing you did

OpenStudy (anonymous):

beware of answer with no splanation

OpenStudy (anonymous):

look

OpenStudy (anonymous):

\[x^2+y^2=12^2\\ 2xx'+2yy'=0\\ x'=-\frac{y}{x}y'\]

OpenStudy (anonymous):

plug in the numbers now

OpenStudy (anonymous):

im getting 4.5929

OpenStudy (anonymous):

by applying the Pythagorean equation we get : \[12^2 = h^2 + w^2\] and when we simplify it to get : \[w=\sqrt{144 - h^2}\] we then differentiate it and plug in the numbers

OpenStudy (anonymous):

@satellite73 i did explained the answer, if you look up you will see a picture of the problem we have